Daytime AO work on 1 June 2011

Observer(s): D. Miller, L Busoni, A. Puglisi, L. Fini, S. Esposito, J Brynnel, C. Kulesa (LBTO)

Telescope Operator: DHuerta (LBTO)

SW Support: (Tucson)

Telescope Support: J. Little (LBTO)

Instrument Support: (LBTO)

Summary

Joar was at the telescope over night

Doug and Craig arrive from Tucson at 1:30

Luca and Alfio stayed at the LBT

Simone and Lorenzo stayed in Safford

Craig installed an old Linux Kernel on the PISCES computer that should be compatible with the PCI card drivers. Testing showed that the interaction was better, but he was still getting a few random timeout because the end of CCD frame read was not detected. A quick test indicates that this new kernel will allow subarray reads. Craig will look into this more this week and confirm this is true. He will make subarrays available via the server once he has tested.

Craig upgraded the PISCES server to now allow a filter, status and non-blocking go commands. Doug tested these commands with his IDL routines and all worked. Craig also installed a cancel command, but this caused some problems so it is not available to use yet. Hopefully in the next couple days cancel will be fixed, all status info will be passed, CCD temperature will be added to status and to header.

After a discussion with Doug, Alfio installed the basic PISCES commands directly into his python routines to collect data during AO engineering.

Craig and Joar left for Tucson at 8:30

-- DougMiller - 02 Jun 2011
Topic revision: r1 - 02 Jun 2011, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback